# Modality Collapse as Mismatched Decoding: Information-Theoretic Limits of Multimodal LLMs

### v1

@@ Abstract
We formalize this as a mismatched decoder problem: a decoder trained on text can only extract information along text-aligned directions. Accessible information is bounded by the Generalized Mutual Information (GMI), with degradation scaling with distributional distance and decoder sensitivity. The bound is a property of the decoder’s scoring rule, not of any particular architecture; it applies whether non-text inputs arrive through a learned projection, a discrete codebook, or no explicit adapter at all. We validate this across five models spanning speech and vision. A controlled experiment (two Prismatic VLMs differing only in encoder text-alignment) confirms the bottleneck is the decoder’s scoring rule, not the encoder or projection. A LoRA intervention demonstrates the fix: training with an emotion objective improves emotion accessibility (++7.5%) without affecting other attributes, confirming that the training objective determines what becomes accessible.
1 Introduction
Multimodal LLMs follow a common recipe: an encoder processes the non-text input (speech, image), a learned projection maps it into the LLM’s embedding space, and the LLM generates text (Liu et al., 2023; Dai et al., 2023; Fixie AI, 2024; Chu et al., 2024; Karamcheti et al., 2024). By standard benchmarks these systems appear to work well: speech models transcribe accurately, vision models caption faithfully. But this success is a mirage. These systems excel at tasks requiring only text-correlated information (the words spoken, the objects visible) while silently discarding everything else. A speech model that transcribes perfectly may have no access to who is speaking or how they feel; a vision model that names every object may be blind to spatial layout or texture. This paper shows that the loss is structural and predictable.
@@ 1 Introduction
Worse, modality-specific information degrades decoding. Removing 64% of the variance in modality-specific directions improves cross-entropy loss; the decoder works better with less information. Lexical probe accuracy nearly doubles from the adapter to the final LLM layer (the decoder amplifies what it knows how to use), while speaker probe accuracy drops by up to 39% over the same path. The mechanism is not active suppression: the decoder simply has no learned response to non-text directions. It is indifferent to them, but that indifference has consequences, because the decoder treats unfamiliar variance as noise. Where the collapse manifests depends on the encoder: non-aligned encoders (Whisper, DINOv2) pass rich representations the decoder is blind to; text-aligned encoders (CLIP, SigLIP) discard non-textual information before it reaches the decoder. Either way, the information is lost (Appendix Table 6).
Our thesis. We formalize this as a mismatched decoder problem (Merhav et al., 1994; Scarlett et al., 2020). The LLM’s scoring rule was learned from text, so it can only extract information along text-aligned directions. Accessible information is bounded by the Generalized Mutual Information (GMI), not the standard mutual information. We call the difference the information accessibility gap. The gap is not about architecture; it is about the scoring rule. No explicit adapter module is required: any system where a text-trained decoder processes non-text representations faces the same constraint, whether the representations arrive through a learned projection, a discrete codebook, or the LLM’s own layers. At inference, every trained model has fixed weights: the scoring rule is whatever the training procedure produced. A model trained predominantly on text has a text-shaped scoring rule regardless of whether the LLM was frozen or fine-tuned during training. What matters is the training objective, not which components were updated (§6.1).
What determines the size of the gap? A linear probe simply measures whether the information it was trained to detect is present. The decoder is a different instrument entirely. It is a deep transformer trained on text, and it is highly sensitive to its input, responding to variance in every direction, including non-text directions (we confirm this empirically via gradient isotropy in §5.5). But its response to non-text variance is unproductive: the decoder interprets unfamiliar structure as noise, and that noise disrupts its text processing. This is why removing modality-specific directions helps: not because the decoder ignores them, but because it responds to them in ways that degrade output. We formalize this: degradation scales with both the distributional distance between modal and text representations and the decoder’s sensitivity to that distance (Theorem 2), and the decoder’s sensitivity is up to $30\times$ that of a probe (Theorem 3, §A.4). Text-aligned encoders reduce the distance (§5.3), but only by discarding non-textual information upstream. The collapse happens earlier, not less.
Contributions. (1) We formalize modality collapse as mismatched decoding and prove that accessible information is bounded by the GMI, with degradation governed by distributional distance and decoder sensitivity (§4). (2) Across five models and two modalities, we demonstrate the information accessibility gap: non-text information survives through the LLM but degrades decoding because the decoder has no learned use for it (§5.2, §5.4). (3) A controlled experiment with Prismatic VLMs (same architecture, same LLM, only encoder text-alignment differs) isolates the scoring rule as the causal variable (§5.3). (4) A LoRA intervention confirms the prescription: training with an emotion objective teaches the decoder to respond to that dimension (++7.5%), while a text-centric objective leaves it indifferent (§6.1).
2 Related work
Multimodal LLMs and the adapter paradigm. The strategy of projecting encoded features into an LLM’s embedding space has become standard for both vision (Liu et al., 2023; Dai et al., 2023; Zhu et al., 2024; Karamcheti et al., 2024; Tong et al., 2024) and speech (Fixie AI, 2024; Chu et al., 2024). These systems achieve strong performance on text-centric benchmarks but struggle with tasks requiring modality-specific information, an observation that has been noted empirically but not explained theoretically.
Modality gap. Liang et al. (2022) showed that CLIP embeddings of different modalities occupy distinct regions of the shared space, separated by a persistent “modality gap.” Huh et al. (2024) observed convergence of representations across modalities. Our work complements these geometric observations with an information-theoretic mechanism: the gap matters not because information is destroyed, but because the decoder cannot access information in unfamiliar directions.
Mismatched decoding. The theory of mismatched decoders (Merhav et al., 1994; Lapidoth, 1996; Scarlett et al., 2020) characterizes the information rate achievable when a decoder is designed for a different channel than the one actually used. We import this framework into the multimodal LLM setting: at inference, any LLM whose scoring rule was shaped by text-dominated training is a mismatched decoder for non-text inputs. The mismatch is determined by the training objective, not by which components were updated.
Representation probing. Linear probes are a standard tool for measuring information content in neural representations (Alain & Bengio, 2017; Conneau et al., 2018; Hewitt & Manning, 2019; Belinkov, 2022). We use probes not to claim information is “encoded” in a representation, but to demonstrate the gap between information present (probe-recoverable) and information accessible (decoder-extractable).
Rate-distortion in ML. Blau & Michaeli (2019) introduced the rate-distortion-perception tradeoff, showing that perceptual quality constraints can increase the rate needed for lossy compression. Our theoretical framework shares the insight that the output distribution shape (here, conformity to the text manifold) imposes additional costs beyond standard rate-distortion.
3 Problem formulation
Having described the phenomenon, we now formalize the architecture and the mismatch that causes it.
Architecture. We consider the standard multimodal LLM pipeline:
$X\xrightarrow{E_{\phi}}H\xrightarrow{P_{\theta}}Z\xrightarrow{L_{\psi}(\cdot,C)}\hat{Y}$ (1)
where $X$ is a non-text input (speech waveform or image), $E_{\phi}$ is a pre-trained encoder (e.g., Whisper (Radford et al., 2023), CLIP (Radford et al., 2021)), $P_{\theta}$ is a learned adapter (linear or MLP projection), $Z\in\mathbb{R}^{d}$ are the hidden states consumed by the LLM $L_{\psi}$, $C$ is the text context (prompt), and $\hat{Y}$ is the generated output. At inference, all components are fixed: the scoring rule $q_{\psi}$ is whatever the training procedure produced. In the common adapter paradigm, only $P_{\theta}$ is trained while $E_{\phi}$ and $L_{\psi}$ are pre-trained; in end-to-end approaches, all components are jointly trained. Either way, at inference the decoder’s scoring rule is fixed, and our theory describes what that fixed scoring rule can extract. The specific form of $P_{\theta}$ varies (linear, MLP, Q-Former (Dai et al., 2023), perceiver, codebook), but the mismatch depends on the scoring rule, not the adapter type (§6.1).
Text law and modal law. Let $q_{\psi}(y|z,c)=L_{\psi}(z,c)$ denote the LLM’s next-token distribution at inference. During pre-training, the LLM processes hidden states drawn from the text law $P_{\!T}(C,Z,Y)$. When processing adapted non-text inputs, the hidden states follow a different modal law $P_{\!M}(C,Z,Y)$.
Assumption 1 (Shared marginal).
The text and modal laws share the same prompts and target tokens: $P_{\!M}(C,Y)=P_{\!T}(C,Y)$.
In other words, the task is the same (e.g., “transcribe this speech”), and only the internal representations $Z$ differ. This isolates the effect of the representation distribution from confounds in the task itself.
Modality collapse. We define modality collapse as the systematic degradation or stagnation of non-text information through a text-trained decoder, depending on the mismatch intensity. For an attribute $A_{\tau}$ of type $\tau$ (e.g., speaker identity, emotion), the information accessibility gap is:
$\Delta_{\text{access}}(\tau)=I(Z;A_{\tau})-\mathrm{GMI}_{P_{\!M}}(A_{\tau}\mid\psi)$ (2)
where $I(Z;A_{\tau})$ is the mutual information (recoverable by an optimal probe) and $\mathrm{GMI}_{P_{\!M}}(A_{\tau}|\psi)$ is the information actually extractable by the decoder’s fixed scoring rule. When $\Delta_{\text{access}}=0$, the decoder extracts everything the probe can; when $\Delta_{\text{access}}$ is large, information is present but inaccessible.
Text-aligned encoders as correlation mappings. Contrastive encoders like CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023) do not merge the modality manifolds (the modality gap persists (Liang et al., 2022)), but they align the informative directions: attributes that co-occur with text descriptions occupy dimensions also used by text representations. When such an encoder feeds an LLM, the modal law $P_{\!M}$ has high overlap with $P_{\!T}$ along informative dimensions, so the distributional distance is small and the accessibility gap is correspondingly small (§A.2 formalizes this via the Wasserstein distance). Non-contrastive encoders like DINOv2 (Oquab et al., 2024) lack this alignment: their informative directions may be orthogonal to $P_{\!T}$, making the information effectively inaccessible to the decoder. We return to this in §5.3, where the Prismatic comparison provides direct evidence.
4 Theoretical framework
The formulation above makes the information loss quantifiable. We establish three results (formal statements, assumptions, and proofs in Appendix A):
1. The decoder has a ceiling. Standard mutual information $I(Z;Y)$ measures how much information $Z$ contains about $Y$ assuming an optimal decoder. But the LLM is not optimal for non-text inputs. The relevant quantity is instead the Generalized Mutual Information (GMI) (Merhav et al., 1994; Scarlett et al., 2020), which measures the maximum rate extractable by a fixed decoder with a given scoring rule. Regardless of how much information a probe can extract from $Z$, the decoder can never exceed $\mathrm{GMI}_{P_{\!M}}(q_{\psi})$ (Theorem 1).
2. The ceiling drops with distributional distance. When modal representations $P_{\!M}$ differ from the text representations $P_{\!T}$ the decoder was trained on, the GMI degrades. We bound this degradation (Theorem 2): it scales with the product of two empirically measurable quantities: (i) the Wasserstein distance $W_{\!1}(P_{\!M},P_{\!T})$, measuring how far modal representations are from text representations, and (ii) the decoder’s Lipschitz constant $L_{\log}$, measuring how sensitive its log-probability is to shifts in the input. The bound makes a clear prediction: models with larger $L_{\log}\cdot W_{\!1}$ should show more degradation, which we validate in §5.5.
[…]
@@ 5 Experimental validation
The bound in Theorem 2 makes three quantities empirically measurable: (i) the Lipschitz constant $L_{\log}$, estimated via gradient norms, measuring the decoder’s sensitivity to input perturbations; (ii) the Wasserstein distance $W_{\!1}(P_{\!M},P_{\!T})$, estimated between modal and text hidden states at matched layers; and (iii) the mode alignment spectrum $\tilde{\alpha}(u_{k})$, which decomposes $W_{\!1}$ geometrically by identifying which directions carry the mismatch. We validate each in turn.
@@ 5.1 Setup
Models. We study five multimodal LLMs spanning two modalities (Table 1). For speech: Ultravox-v0.6 (Fixie AI, 2024) (Whisper encoder, MLP adapter with SwiGLU, Llama-3.1-8B) and Qwen2-Audio-7B (Chu et al., 2024) (Whisper encoder, linear adapter, Qwen2-7B). For vision: LLaVA-v1.5-7B (Liu et al., 2023) (CLIP-ViT, MLP adapter, Vicuna-7B) and two Prismatic VLMs (Karamcheti et al., 2024) sharing the same MLP adapter and Vicuna-7B backbone but differing only in their vision encoder: DINOv2 (not text-aligned) and SigLIP (text-aligned via contrastive loss).
@@
Probing protocol. We extract representations at four hook points per model: encoder output, adapter output, LLM layer 16, and LLM final layer (after LayerNorm). Representations are mean-pooled across sequence positions (restricting to audio-only tokens yields nearly identical results; Appendix Table 9). We train $\ell_{2}$-regularized logistic regression probes (5 seeds, 80/20 stratified split, $z$-score normalization) for each combination of hook point, information type, and dataset.
@@
The Prismatic pair (shared architecture, different encoder) and the two speech models (shared encoder, different LLM) provide controlled contrasts that isolate specific variables.

### v2

@@ Abstract
Numerous studies have shown that multimodal LLMs process speech and images well but fail in non-intuitive ways rendering trivial tasks such as object counting unreliable. We investigate this behavior from an information-theoretic perspective by framing multimodal LLM inference as a mismatched decoder problem: a decoder trained primarily on text can only extract information along text-aligned directions (removing up to 98% of the variation in modality-specific (non-text) directions improves decoder loss) and the amount of accessible information is bounded by the Generalized Mutual Information (GMI). We show that information loss is bounded as the distributional mismatch between the source data and the text data increases, and as the sensitivity of the decoder increases. This bound is a function of the model’s scoring rule not its architecture. We validate the predictions across five models spanning speech and vision. A controlled study (two Prismatic VLMs differing only in encoder text-alignment) shows that the bottleneck lies in the scoring rule of the decoder rather than the text-alignment of the encoder or the learned projection. A LoRA intervention demonstrates that simply training with an emotion-related objective improves emotion detection from 17.3% to 61.8% task accuracy without affecting other attributes, confirming that the training objective determines what becomes accessible.
1 Introduction
Multimodal LLMs follow a common recipe: an encoder processes the non-text input (speech, image), a learned projection maps it into the LLM’s embedding space, and the LLM generates a text response (Liu et al., 2023; Dai et al., 2023; Fixie AI, 2024; Chu et al., 2024; Karamcheti et al., 2024). These systems appear to perform well on standard benchmarks. However, there are many reported gaps in model performance (Chen et al., 2026; Cuervo et al., 2026; Thrush et al., 2022; Yüksekgönül et al., 2023; Jain et al., 2025) where even some trivial tasks are not resolved. A model might produce a perfect response to speech input but have no understanding of the emotional state of the speaker; it may correctly identify every object in an image and yet fail to count the number of objects or understand the spatial relationships. Here, the information is clearly available to the model, since it can list every object, but it cannot make use of this available information for all tasks. We refer to this selective failure as modality collapse.
To provide insight into this behavior we borrow the well-defined concept of a mismatched decoder from communication theory, where a decoder originally designed for source A receives source B. A multimodal LLM fits this definition of a mismatched decoder exactly: an LLM trained to ”decode” text representation is now expected to extract information from audio and/or image projections. Our mismatched decoder analysis is a formalization of the perception constraint cost (Blau and Michaeli, 2019).
In this framing the accessible information is bounded by the Generalized Mutual Information (GMI) not standard mutual information, as would be the case for matched decoders. We refer to this difference as the information accessibility gap. The gap is not about architecture; it is about the scoring rule. No explicit adapter module is required: any system where a text-trained decoder processes non-text representations faces the same constraint, whether the representations arrive through a learned projection, a discrete codebook, or the LLM’s own layers. At inference, every trained model has fixed weights: the scoring rule is whatever the training procedure produced. A model trained predominantly on text has a text-shaped scoring rule regardless of whether the LLM was frozen or fine-tuned during training. What matters is the training objective, not which components were updated (§6.1).
What causes this gap? A transformer model is sensitive to its input in all directions; one trained on text is still sensitive to non-text directions (we validate this empirically using gradient isotropy in §5.6). That said, its sensitivity to non-text directions is unproductive. The decoder treats unfamiliar structure as noise, and noise hurts its text processing, as would be expected in a mismatched decoder. In §5.4 we empirically show that removing up to 98% of the variation in modality-specific (non-text) directions improves decoder loss; the decoder is not indifferent to non-text directions, but responds to them in destructive ways. We can quantify this degradation (Theorem 2). It is proportional to both the distributional distance between the modality and text representations and the decoder’s sensitivity to this distance (its Lipschitz constant $L_{\log}$, which measures how much the output changes per unit change in the input). This sensitivity can be up to $30\times$ that of a linear probe (a classifier trained on frozen representations to test what information is present; Theorem 3, §A.5). Text-aligned encoders reduce the distance (§5.5), but only by discarding non-textual information earlier on. The collapse occurs earlier, not less.
Contributions. (1) We formalize modality collapse as mismatched decoding, and prove that the accessible information is at most the GMI, and degrades with the distributional mismatch and the sensitivity of the decoder (§4). (2) On five models and two modalities, we show that the information accessibility gap is present: non-text information is preserved by the LLM but is not decodable because the decoder is not incentivized to use it (§5.4, §5.2). (3) We run a controlled study on Prismatic VLMs (identical architecture, identical LLM, only text-alignment of the encoder changes), and show that the scoring rule is the causal factor (§5.5). (4) We run a LoRA intervention that confirms the prescription: training with an emotion objective improves emotion task accuracy from 17.3% to 61.8% (++7.5% probe), while training with a text objective does not (§6.1).
2 Related work
Multimodal LLMs and the adapter paradigm. The practice of projecting encoded representations into the embedding space of an LLM has been widely adopted for both vision (Liu et al., 2023; Dai et al., 2023; Zhu et al., 2024; Karamcheti et al., 2024; Tong et al., 2024) and speech (Fixie AI, 2024; Chu et al., 2024). While these approaches perform well on text-centric tasks, they perform poorly on tasks that require modality-specific knowledge, a phenomenon observed empirically but not understood theoretically.
Modality gap. Liang et al. (2022) observed that the CLIP embeddings of different modalities are separated by a “modality gap” in the shared space. More recently, Huh et al. (2024) reported that the representations of different modalities become similar. While our results are largely orthogonal to these geometric findings, we provide an information-theoretic explanation for the modality gap: it is not that information is lost, but rather that the decoder is unable to make use of information in unfamiliar directions.
Modality collapse. Javaloy et al. (2022) use this term to describe models that discard an entire modality during training due to gradient conflicts. Our usage describes a subtler failure: the model processes all modalities and preserves their information internally, but the decoder selectively fails to use non-primary-modality content.
Mismatched decoding. The mismatched decoder framework (Merhav et al., 1994; Lapidoth, 1996; Scarlett et al., 2020) provides the data rates supported by a decoder mismatched to its incoming channel. We borrow this approach and apply it to the multimodal LLM framework. At inference time, any LLM that uses a scoring rule that was learned via a text-dominant training procedure is a mismatched decoder when applied to non-text inputs. The mismatch arises from the loss function, not from what parameters were updated.
Representation probing. Linear probes are a common method for probing the information content of neural representations (Alain and Bengio, 2017; Conneau et al., 2018; Hewitt and Manning, 2019; Belinkov, 2022). We do not use probing to argue that some information is “encoded” in a representation, but to highlight the difference between information that is present (probe-recoverable) and information that is accessible (decoder-extractable).
Rate-distortion-perception (RDP) in ML. Blau and Michaeli (2019) formalized perceptual quality as a distributional constraint: reconstructions are perceptually good when their distribution matches the real data distribution. The text-trained decoder imposes a perception constraint (text-aligned directions), forcing the adapter to compromise modality-specific information to satisfy the constraint. We draw on this framework in §3, where the decoder takes the role of the perceptual evaluator.
3 Problem formulation
Architecture. We consider the typical multimodal LLM pipeline:
$X\xrightarrow{E}H\xrightarrow{A}Z\xrightarrow{L(\cdot,C)}\hat{Y}$ (1)
where $X$ is a non-text input (speech waveform or image), $E$ is a pre-trained encoder (e.g., Whisper (Radford et al., 2023), CLIP (Radford et al., 2021)), $A$ is a learned adapter, $Z\in\mathbb{R}^{d}$ are the hidden states consumed by the LLM $L$, $C$ is the text context (prompt), and $\hat{Y}$ is the generated output. The specific form of $A$ varies (linear, MLP, Q-Former (Dai et al., 2023), perceiver, codebook), but the mismatch depends on the scoring rule, not the adapter type (§6.1).
Scoring rule, text law and modal law. Let $q(y|z,c)$ denote the LLM’s next-token distribution at inference. This is the fixed scoring rule resulting from training. During pre-training, the LLM processes hidden states drawn from the text law $P_{\!T}(C,Z,Y)$. When processing adapted non-text inputs, the hidden states follow a different modal law $P_{\!M}(C,Z,Y)$.
We assume the text and modal laws share the same prompts and target tokens ($P_{\!M}(C,Y)=P_{\!T}(C,Y)$), so the only difference is in the representations $Z$. The quantitative bound (§4) additionally requires that the decoder’s log-score is Lipschitz continuous and that representations occupy a bounded region (Assumptions 23 in Appendix A.2, validated empirically in §5.6).
Formalizing modality collapse. The earlier intuitive explanation of modality collapse (information is available but the decoder cannot use it) can be stated precisely. For an attribute $S_{\tau}$ of type $\tau$ (e.g., speaker identity, emotion), we define the information accessibility gap as:
$\Delta_{\text{access}}(\tau)=I(Z;S_{\tau})-\mathrm{GMI}_{P_{\!M}}(S_{\tau}\mid q)$ (2)
where $I(Z;S_{\tau})$ is the information about $S_{\tau}$ present in the representation and $\mathrm{GMI}_{P_{\!M}}(S_{\tau}|q)$ is the information actually extractable by the decoder’s fixed scoring rule. When $\Delta_{\text{access}}=0$, the decoder extracts everything present; when $\Delta_{\text{access}}$ is large, information is present but inaccessible.
@@ 3 Problem formulation
Text-aligned encoders as correlation mappings. Contrastive encoders like CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023) do not merge the modality manifolds (the modality gap persists (Liang et al., 2022)), but they align the informative directions: attributes that co-occur with text descriptions occupy dimensions also used by text representations. When such an encoder feeds an LLM, the modal law $P_{\!M}$ has high overlap with $P_{\!T}$ along informative dimensions, so the distributional distance is small and the accessibility gap is correspondingly small (we formalize this in terms of the Wasserstein distance in §A.2). Non-contrastive encoders like DINOv2 (Oquab et al., 2024) lack this alignment: their informative directions may be orthogonal to $P_{\!T}$, making the information effectively inaccessible to the decoder. We return to this in §5.5, where the Prismatic comparison provides direct evidence.
4 Theoretical framework
The mismatched decoder framing makes three concrete predictions. First, since the decoder was trained on text, there should be an upper limit on how much non-text information it can extract, regardless of how much is present in the representation. Second, this limit should get worse as the modality representations diverge from text. Third, a simple classifier (a linear probe) trained directly on the representations should not be subject to the same limit, since it is not constrained by a text-trained scoring rule. We formalize each below (formal statements, assumptions, and proofs in Appendix A):
1. The decoder has an upper bound. The standard mutual information $I(Z;Y)$ represents the information $Z$ contains about $Y$ assuming an optimal decoder. But the LLM is not an optimal decoder for non-text inputs, given a text-focused training objective. Instead, the relevant quantity is the Generalized Mutual Information (GMI) (Merhav et al., 1994; Scarlett et al., 2020), the maximum rate achievable with a fixed scoring rule. No matter how much information any probe can get from $Z$, the decoder is always bounded by $\mathrm{GMI}_{P_{\!M}}(q)$ (Theorem 1).
2. The ceiling drops with distributional distance. The GMI decreases when the modal representations $P_{\!M}$ differ from the text representations $P_{\!T}$ the decoder was trained on. How much it drops depends on two measurable quantities: how far the modal representations are from text (the Wasserstein distance $W_{\!1}$), and how sensitive the decoder is to that difference (the Lipschitz constant $L_{\log}$). The drop is bounded by their product (Theorem 2). Both quantities can be estimated from a trained model ($W_{\!1}$ from the representations, $L_{\log}$ from gradient norms), so the bound yields a testable prediction: models with larger $L_{\log}\cdot W_{\!1}$ should show more degradation. We validate this ranking in §5.6. In the high-mismatch regime the bound is numerically loose, but it remains an ordinal diagnostic: it correctly ranks model–modality pairs by collapse severity and predicts which information types degrade (Appendix A.3 tightens the prefactor from ${\sim}e^{150}$ to ${\sim}e^{7}$).
@@ 4 Theoretical framework
The three propositions above establish the perception-distortion face of the RDP tradeoff: as the distributional distance $W_{\!1}(P_{\!M},P_{\!T})$ grows, accessible information (GMI) decreases, and the decoder is disproportionately more affected than a probe.
@@ 5.1 Setup
Models. We investigate five multimodal LLMs across two modalities (Table 1). For speech: Ultravox-v0.6 (Fixie AI, 2024) (Whisper encoder, MLP adapter with SwiGLU, Llama-3.1-8B) and Qwen2-Audio-7B (Chu et al., 2024) (Whisper encoder, linear adapter, Qwen2-7B). For vision: LLaVA-v1.5-7B (Liu et al., 2023) (CLIP-ViT, MLP adapter, Vicuna-7B) and two Prismatic VLMs (Karamcheti et al., 2024) with the same MLP adapter and Vicuna-7B backbone but with different vision encoders: DINOv2 (not text-aligned) and SigLIP (text-aligned via contrastive loss).
@@
Probing protocol. We consider four hook points per model: encoder output, adapter output, LLM layer 16, and LLM final layer (after LayerNorm). Representations are mean-pooled across sequence positions. Ultravox also provides audio token boundaries, but pooling over audio-only tokens results in nearly identical trends (see Appendix Table 9). For each hook point, information type and dataset combination, we train $\ell_{2}$-regularized logistic regression probes (5 seeds, 80/20 stratified split, $z$-score normalization).
@@
The Prismatic pair (shared architecture, different encoder) and the two speech models (shared encoder, different LLM) serve as controls that allow us to isolate these specific variables.
5.2 The information accessibility gap in practice
Linear probes across layers show that for all 5 models, non-text information is not destroyed by the LLM (Appendix Tables 5 and 7): at the final layer, speaker identity is $22\times$ above chance for Ultravox and object categories are $55\times$ above chance in Prismatic-D (.792.792 vs. .014.014). The information is preserved. Does the decoder use it?
The answer depends on text-alignment. In both speech models, the LLM approximately doubles lexical information (++92% Ultravox, ++95% Qwen2-Audio), but degrades speaker identity (-8% to -39%). When the encoder is text-aligned (CLIP in LLaVA, SigLIP in Prismatic-S), $P_{\!M}\approx P_{\!T}$ and all information types improve (++5–8%). The decoder amplifies what it was trained on and is indifferent to the rest. The LISTEN benchmark (Chen et al., 2026) independently observed this pattern: audio LLMs predict emotions based on what the speaker says rather than how they say it. Our framework explains why: the text-shaped scoring rule amplifies lexical content (++92–95%) while remaining indifferent to acoustic features, so emotion predictions are driven by word choice rather than prosody.
The same pattern appears in vision, though in a weaker mismatch regime (Appendix Table 10). Three visual features not present in text (object count, object size, spatial distribution) show flat or modest improvement through the LLM ($L_{\log}\cdot W_{\!1}\approx 13$$54$ vs. $162$ for speech), consistent with stagnation unlike the sharp degradation in speech.
5.3 Mode alignment
The bound describes how much information is lost; mode alignment tells us the directions in which it is lost. We perform a principal component analysis of the adapter’s output representations under the modal distribution, yielding a covariance matrix $\Sigma_{M}$ with eigenvectors $u_{k}$ and eigenvalues $\lambda_{k}$, sorted by decreasing eigenvalue. We refer to each eigenvector as a mode: Mode 0 is the direction along which the adapter’s output varies the most, Mode 1 the next, and so on. For each mode, we ask: has the decoder seen variation along this direction during training? The alignment score
$\rho_{k}=u_{k}^{\top}\Sigma_{T}\,u_{k}\,/\,\lambda_{k}$
is the Rayleigh quotient of the text covariance $\Sigma_{T}$ evaluated along the modal eigenvector $u_{k}$, normalized by the modal eigenvalue $\lambda_{k}$. The numerator measures how much the text distribution varies along direction $u_{k}$; the denominator is how much the modality varies along the same direction. When $\rho_{k}=1$, the text and modal distributions have equal variance along $u_{k}$, so the decoder has been trained on exactly this kind of variation and can use it. When $\rho_{k}\approx 0$, the modality varies along a direction where text has almost no variance: the decoder has never seen input spread along this direction and cannot productively respond to it. We label modes with $\rho_{k}<0.5$ as modality-specific (MS); the threshold is a practical cutoff, and results are not sensitive to the exact value (Appendix Table 11).
For non-aligned speech encoders, Mode 0 (the direction of largest adapter variance) is modality-specific ($\rho_{0}\leq 0.034$), accounting for 51–79% of adapter variance. For Prismatic-DINOv2, Mode 0 is also modality-specific ($\rho_{0}=0.034$) but carries only 8% of variance; the MS variance is instead distributed across 53 modes (71% total; Table 11). In both cases, the dominant source of variation in the adapter’s output is invisible to the text distribution. For SigLIP, even Mode 0 is text-aligned ($\rho_{0}=0.83$): the encoder has already organized its output along text-like directions. As representations pass through the LLM, text-aligned variance is amplified (${\sim}150\times$) while MS variance is unaffected: MS modes go from 63.6% of adapter variance to <<1% at the final layer in Ultravox, not destroyed but drowned out.